OpenStudy (anonymous):

a body dropped from a tower travels half of the distance in the last second of its motion the total time of fall will be?(g=10m/s^2)

OpenStudy (anonymous):

do you know the formula for the distance travelled in nth second of motion?

OpenStudy (anonymous):

yes d=u+a(n-1/2)

OpenStudy (ajprincess):

d/2=10(n-1/2) d=10(2n-1) d=1/2*10*n^2 =5n^2 5n^2=20n-10 5n^2-20n+10=0 n^2-4n+2=0 n^2-4n+4=2 (n-2)^2=2 n-2=sqrt2 n=sqrt2+2

OpenStudy (anonymous):

s u r correct